[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Nov 18, 2015
Christopher Mello - NOAA Federal
christopher.mello at noaa.gov
Wed Nov 18 13:28:38 UTC 2015
Here is today's update on NHC's computer operations.
*TAFB will backup HFO on Thursday. *
*TSB will finish the AWIPS2 EDEX hardware refresh this morning. *
A few issues popped up during the AWIPS2 hardware refresh. The AWIPS2
text editor now requires the full pil to call and send out products in the
AWIPS2 text editor. Chris will investigate.
Matt researched the issues regarding SFC analysis files failing to
transfer between OPC and NHC. Besides a couple of random FAILEDs in the
dbnet_launch.pl log, things appeared normal . Both self-corrected on
the next cron run, so these were just minor, random hiccups.
The dbnet_launch.pl script sends the 'pre' TAFB surface analysis VGFs to
our DBNet server and then on to NCO. I checked the DBNet logs, and didn't
see any obvious problems there, either.
If this continues, we need to coordinate with NCO/OPC to determine if
networks issues are causing failures.
Forecasters reported issues with the word rap in the AWIPS2 text editor
since last weeks upgrade to 15.1.1 Chris will investigate.
No new issues.
National Hurricane Center
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts